# Learning Knowledge-based Prompts for Robust 3D Mask Presentation Attack Detection

### v1

@@ 3.3 Overall Objective and Inference
The overall optimization objective is to minimize , where is a hyperparameter. During the inference phase, in Equation 10 is employed to evaluate the similarity of the image feature and the knowledge-based text feature of category . Then, all the similarity values are aggregated as follows:
This set of similarity scores is utilized as the predictive probability distribution of the trained model over the label set. Subsequently, a conventional threshold-based evaluation method is applied to determine whether the test sample is a real face or a 3D mask.
@@ 4.1 Experimental Setups
Datasets. We evaluate the performance of the proposed method on three widely used datasets: 3DMAD (3D Mask Attack Dataset) [40], HKBU-MARs V1+ (Hong Kong Baptist University 3D Mask Attack with Real-World Variations) [33] and HiFiMask (High-Fidelity Mask dataset) [11]. Figure 6 shows examples from three datasets. Note that only RGB images are used in our experiments.
@@
We use the widely-used metrics to evaluate the proposed method against the state-of-the-art approaches, including Half Total Error Rate (HTER), Equal Error Rate (ERR), Area Under Curve (AUC), Attack Presentation Classification Error Rate (APCER), and Bonafide Presentation Classification Error Rate (BPCER). Here, “B@A=0.1/0.01” represents the BPCER when APCER equals 0.1 or 0.01. Note that EER is calculated based on the development set, which shares a similar data distribution with the training set. In contrast, HTER is computed using thresholds derived from the development set but applied to a test set with significantly different data distributions. Hence, EER reflects the model’s ability to fit the training data, while HTER measures its generalization capability on unseen test data.
Competitors. We select nine representative face presentation attack detection approaches for performance comparison. They are listed as follows. 1) MS-LBP [52] and CTA [53]: Classical methods that extract color and texture features using handcrafted feature descriptors. 2) CDCN++ [54]: A widely used approach that leverages central difference convolutional neural networks to learn fine-grained features. 3) HRFP [10]: A method that extracts fine-grained information from local image patches and was the winner of the 3D High-Fidelity Mask Attack Detection Challenge at ICCV 2021. 4) ViTranZFAS [55] and MD-FAS [56]: Approaches that utilize pre-trained models to differentiate between real and fake faces. 5) CFrPPG [57] and LeTSrPPG [58]: Techniques that learn remote photoplethysmography (rPPG) features, demonstrating strong discriminative capability for 3D mask presentation attack detection. 6) FASTEN [13]: A recent state-of-the-art method that learns spatiotemporal features, achieving superior performance on commonly used datasets. To ensure fairness, we follow the settings of the original papers and use the performance data reported in those studies for comparison. If the original papers do not report performance data for certain protocols, we mark them with ”U”, indicating that the data is unavailable.
@@ 4.2.1 Intra-Dataset Evaluation
Compared to FASTEN, the proposed method achieves HTER improvements of 71.79%, 22.54%, and 27.96% on the 3DMAD, HKBU-MARs V1+, and HiFiMask datasets, respectively. The proposed method also achieves significant performance gains across the other four evaluation metrics: EER, AUC, B@A=0.1, and B@A=0.01. Notably, while FASTEN requires multiple frames per video for 3D mask presentation attack detection, the proposed method requires only one single frame each video, yet achieves significant performance gains. These results demonstrate that integrating knowledge graphs and causal prompt learning effectively adapts pre-trained vision-language models for 3D mask presentation attack detection.
@@ 4.2.2 Cross-Dataset Evaluation
Comparison of feature visualizations. We use t-SNE to visualize the image feature space of the target domain learned by FASTEN and the proposed method under the HiFiMask3DMAD protocol, as shown in Figure 7. Since FASTEN uses multi-frame timing data, each point in Figure 7 represents one video. For both 3D masks and real faces, the sample distributions mapped by FASTEN are relatively scattered within each category. In contrast, the proposed method shows more compact clustering within the same category, reflecting greater robustness to variations in unseen scenarios. Furthermore, the inter-class distance between different categories in the feature space learned by the proposed method is larger than FASTEN, demonstrating a significantly larger separation between classes. Notably, as show in Figure 7(b), our method aggregates real faces into two clusters. In the 3DMAD dataset, the real face data originate from two distinct scenarios. Our approach tends to map samples with similar category-irrelevant factors together.
@@ 4.3 Comparison with Classical Prompt Learning Methods
| HTER | EER | AUC | B@A=0.1 | B@A=0.01 | |
| 18.10 | 5.53 | 99.75 | 0.71 | 1.03 | |
| CoOp [21] | 12.39 | 4.55 | 99.66 | 0.77 | 1.09 |
| FLIP [23] | 4.77 | 1.79 | 99.83 | 0.80 | 1.04 |
| Ours | 1.04 | 1.76 | 99.97 | 0.05 | 0.56 |
Comparison of feature visualizations. We use t-SNE to visualize the image feature space of CoOp and the proposed method, with results shown in Figure 9. For this visualization, we randomly select 1,000 samples each from the training and test sets. Each point in Figure 9 represents an individual image.
The HiFiMask dataset contains highly realistic 3D masks, leading CoOp to cluster 3D masks and real faces from the HiFiMask dataset closely together. This indicates that CoOp’s image encoder, without fine-tuning, struggles to adaptively distinguish between real faces and highly realistic 3D masks. Consequently, the classification boundaries and thresholds established on the training set fail to generalize effectively to the test set.
@@ 4.4 Visualization and Analysis
4.4.1 Analysis of the Correlation between Knowledge Graph Entity and Images
Knowledge graph entities play a pivotal role in knowledge-based prompt learning. Analyzing the effect of different entities in the learning process is beneficial for the dissection of our method. We visualize the attention scores between entities and images in the visual-specific knowledge filter to achieve this. The results under the HiFiMask3DMAD protocol are shown in Figure 10. For 3D masks, entities related to material properties, such as “resin” and “plastic”, structure features like “3D” and “edge” (e.g., between the mask and the face), and associated terms, including “spoof face”, “face mask”, and “evil” receive significant attention during the decision-making process. In contrast, for real faces, entities focusing on material characteristics like “skin”, texture details such as “natural details” and “natural face”, and actions like “blinking” and “subtle movement”, as well as terms like “genuine face”, “legal”, and “verified” are highly attended. These observations align closely with commonly identified detection cues in 3D mask presentation attack detection. Notably, numerous attended entities represent semantic extensions of class names tailored to the business context of 3D mask detection. Simple class labels such as 3D masks and real faces fail to encapsulate the nuanced information required for robust detection. By explicitly incorporating prior knowledge, the proposed framework generates accurate and comprehensive prompts aligned with practical detection scenarios, significantly improving model performance.
@@ 4.4.2 Analysis of Spurious Correlation Elimination
The spurious correlation elimination learning paradigm enables the detection model to better focus on generic yet category-relevant local patches. We perform a visualization analysis of category-relevant and category-irrelevant local patches involved in spurious correlation elimination learning to validate this phenomenon. The results under the HiFiMask3DMAD protocol are illustrated in Figure 11. We train the model using the HiFiMask dataset and then visualize the local patches calculated by Equation 7 within the 3DMAD dataset.
The distribution patterns of category-relevant local patches differ significantly between real faces and 3D masks. For a given category, these patches exhibit a relatively consistent distribution across different individuals. In real faces, the patches are concentrated in areas rich in three-dimensional structures, such as the nose and eyes. In contrast, 3D masks focus on the covered mask areas and the junctions between the mask and the face. These substantial inter-category differences demonstrate the model’s strong discriminative capabilities. Furthermore, the consistent discriminative cues within each category underscore the model’s good generalization ability.
@@ 4.5.4 Robustness Analysis of the Number of Entities
Entities play a pivotal role in knowledge-based prompt learning. We further analyze the impact of entity quantity on the performance of 3D mask presentation attack detection. Figure 13 shows the model’s performance after randomly removing one entity from the knowledge graph under the HiFiMask3DMAD protocol, repeated ten times. Notably, when an entity is removed, the descriptions corresponding to the triples associated with that entity are also excluded.
Although removing different entities leads to slight performance declines, the overall fluctuation across the five metrics remains minimal. That is, the proposed method is robust to the removal of a single entity. As the proposed model utilizes fine-grained discriminative descriptions to extract prior knowledge from large-scale language models, it can effectively expand the entities. The information loss caused by removing a single entity is likely compensated by the expansion of other related entities, resulting in negligible performance variations.
To further assess the model performance to varying entity quantities, we also conduct a differential T-test on the data distribution and variance across the four groups under the HiFiMask3DMAD protocol (i.e., randomly removing one, three, six, and ten entities from the knowledge graph). Figure 14 illustrates the results, where “ns” indicates no significant difference, and additional asterisks (*) denote higher significance levels. Overall, as more entities are removed, the model’s performance deteriorates. The degradation occurs because larger removals decrease the likelihood of compensating for lost information through other entities. The results demonstrate the critical role of knowledge graph entities in enhancing 3D mask presentation attack detection performance.

### v2

@@ 3.3 Overall Objective and Inference
The overall training objective combines the two proposed regularizations and is defined as:
where is a hyperparameter that balances the contribution of the two regularizations.
During inference, we compute the predicted probability distribution of a given sample based on Equation 10. Specifically, the trained model is first used to compute for each category according to Equation 10, obtaining the similarity scores between the image feature and the knowledge-based text feature of the -th category. All the similarity scores are aggregated to form a set as follows:
This set of similarity scores serves as the predicted probability distribution over the label set. A threshold-based decision rule is subsequently applied to classify the test sample as either a real face or a 3D mask. If the predicted probability of a test sample being a real face exceeds the threshold, the sample is classified as a real face; otherwise, it is classified as a 3D mask. The decision threshold is determined on the development set and corresponds to the point at which the False Rejection Rate (FRR) equals the False Acceptance Rate (FAR), ensuring a balanced trade-off between the two types of classification errors.
@@ 4 Experiments
This section presents comprehensive experimental evaluations of the proposed method. We begin by describing the experimental setup, including datasets, protocols, and implementation details. Next, we compare the proposed approach with existing state-of-the-art methods under both intra-dataset and cross-dataset evaluation settings. Finally, we perform extensive ablation studies and provide qualitative visualizations to analyze the effectiveness of each component within our framework.
@@ 4.1 Experimental Setups
Datasets. We evaluate the performance of the proposed method on three widely used datasets: 3DMAD (3D Mask Attack Dataset) [39], HKBU-MARs V1+ (Hong Kong Baptist University 3D Mask Attack with Real-World Variations) [32] and HiFiMask (High-Fidelity Mask dataset) [10]. Note that only RGB images are used in our experiments.
@@
We use the widely-used metrics to evaluate the proposed method against the state-of-the-art approaches, including Half Total Error Rate (HTER), Equal Error Rate (EER), Area Under Curve (AUC) [50], Attack Presentation Classification Error Rate (APCER), and Bonafide Presentation Classification Error Rate (BPCER) [51]. Here, “B@A=0.1/0.01” represents the BPCER when APCER equals 0.1 or 0.01. Note that EER is calculated based on the development set, which shares a similar data distribution with the training set. In contrast, HTER is computed using thresholds derived from the development set but applied to a test set with significantly different data distributions. Hence, EER reflects the model’s ability to fit the training data, while HTER measures its generalization capability on unseen test data.
Competitors. We select nine representative face presentation attack detection approaches for performance comparison. They are listed as follows. 1) MS-LBP [55] and CTA [56]: Classical methods that extract color and texture features using handcrafted feature descriptors. 2) CDCN++ [57]: A widely used approach that leverages central difference convolutional neural networks to learn fine-grained features. 3) HRFP [9]: A method that extracts fine-grained information from local image patches and is the winner of the 3D High-Fidelity Mask Attack Detection Challenge at ICCV 2021. 4) ViTranZFAS [58] and MD-FAS [59]: Approaches that utilize pre-trained models to differentiate between real and fake faces. 5) CFrPPG [60] and LeTSrPPG [61]: Techniques that learn remote photoplethysmography (rPPG) features, demonstrating strong discriminative capability for 3D mask presentation attack detection. 6) FASTEN [12]: A recent state-of-the-art method that learns spatiotemporal features, achieving superior performance on commonly used datasets. To ensure fairness, we follow the settings of the original papers and use the performance data reported in those studies for comparison. If the original papers do not report performance data for certain protocols, we mark them with ”U”, indicating that the data is unavailable.
@@ 4.2 Comparison with State-of-the-Art Methods
In this subsection, we compare and analyze our method against existing state-of-the-art methods under both intra-dataset and cross-dataset protocols.
@@ 4.2.1 Intra-Dataset Evaluation
Compared to FASTEN, the proposed method achieves HTER improvements of 71.79%, 22.54%, and 27.96% on the 3DMAD, HKBU-MARs V1+, and HiFiMask datasets, respectively. The proposed method also achieves significant performance gains across the other four evaluation metrics: EER, AUC, B@A=0.1, and B@A=0.01. Notably, while FASTEN requires multiple frames as input for 3D mask presentation attack detection, the proposed method requires only one single frame as input, yet achieves significant performance gains. These results demonstrate that integrating knowledge graphs and causal prompt learning effectively adapts pre-trained vision-language models for 3D mask presentation attack detection.
@@ 4.2.2 Cross-Dataset Evaluation
Comparison of feature visualizations. We employ t-SNE to visualize the image feature distributions of the test domain, as extracted by FASTEN and our method under the HiFiMask3DMAD protocol, as shown in Figure 8. Since FASTEN utilizes multi-frame temporal information, each point in Figure 8 corresponds to one video sample. It can be observed that the sample features extracted by FASTEN exhibit dispersed intra-class distributions and narrow inter-class margins. In contrast, our method produces sample features with compact intra-class distributions and significantly wider inter-class separations. Since the test domain is not accessible during training, the distribution of features indicates the generalization capability of the learned models. Although the visualization of FASTEN suggests that the test samples appear to be linearly separable, the dispersed feature layout and narrow inter-class margins imply that the decision boundary derived from the training data may not perfectly traverse this narrow margin, potentially leading to misclassifications. In contrast, the wider inter-class margins and tighter intra-class clustering produced by our method enable the decision boundary to generalize more robustly to the test domain, as it is more likely to fall within the expanded inter-class margin. Notably, as show in Figure 8(b), our method aggregates real faces into two clusters. In the 3DMAD dataset, the real face data originate from two distinct scenarios. Our approach tends to map samples with similar category-irrelevant factors together.
@@ 4.3 Comparison with Classical Prompt Learning Methods
Comparison of feature visualizations. We use t-SNE to visualize the image feature space of CoOp and the proposed method, with results shown in Figure 10. For this visualization, we randomly select 1,000 samples each from the training and test sets. Each point in Figure 10 represents an individual image. The HiFiMask dataset contains highly realistic 3D masks, leading CoOp to cluster 3D masks and real faces from the HiFiMask dataset closely together. This indicates that CoOp’s image encoder, without fine-tuning, struggles to adaptively distinguish between real faces and highly realistic 3D masks. Consequently, the classification boundaries and thresholds established on the training set fail to generalize effectively to the test set.
@@ 4.4 Visualization and Analysis
We perform a visual analysis of the correlations between knowledge graph entities and images, as well as category-relevant and category-irrelevant patches within images.
4.4.1 Analysis of the Correlations between Knowledge Graph Entities and Images
Knowledge graph entities play a pivotal role in knowledge-based prompt learning. Analyzing the effect of different entities in the learning process is beneficial for the dissection of our method. We visualize the attention scores between entities and images in the visual-specific knowledge filter to achieve this. The results under the HiFiMask3DMAD protocol are shown in Figure 11. For 3D masks, entities related to material properties, such as “resin” and “plastic”, structure features like “3D” and “edge” (e.g., between the mask and the face), and associated terms, including “spoof face”, “face mask”, and “evil” receive significant attention during the decision-making process. In contrast, for real faces, entities focusing on material characteristics like “skin”, texture details such as “natural details” and “natural face”, and actions like “blinking” and “subtle movement”, as well as terms like “genuine face”, “legal”, and “verified” are highly attended. These observations align closely with commonly identified detection cues in 3D mask presentation attack detection. Notably, numerous attended entities represent semantic extensions of class names tailored to the business context of 3D mask detection. Simple class labels such as 3D masks and real faces fail to encapsulate the nuanced information required for robust detection. By explicitly incorporating prior knowledge, our framework generates accurate and comprehensive prompts aligned with practical detection scenarios, significantly improving model performance.
@@ 4.4.2 Analysis of Spurious Correlation Elimination
The spurious correlation elimination learning paradigm enables the detection model to better focus on generic yet category-relevant local patches. We perform a visualization analysis of category-relevant and category-irrelevant local patches involved in spurious correlation elimination learning to validate this phenomenon. The results under the HiFiMask3DMAD protocol are illustrated in Figure 12. We train the model using the HiFiMask dataset and then visualize the local patches calculated by Equation 7 within the 3DMAD dataset. The distribution patterns of category-relevant local patches differ significantly between real faces and 3D masks. For a given category, these patches exhibit a relatively consistent distribution across different individuals. In real faces, the patches are concentrated in areas rich in three-dimensional structures, such as the nose and eyes. In contrast, 3D masks focus on the covered mask areas and the junctions between the mask and the face. These substantial inter-category differences demonstrate the model’s strong discriminative capabilities. Furthermore, the consistent discriminative cues within each category underscore the model’s good generalization ability.
@@ 4.5.3 Analysis of the Hyperparameter for Spurious Correlation Elimination Regularization
| HTER | EER | AUC | B@A=0.1 | B@A=0.01 | |
| CLIP ViT-B/16 | 1.04 | 1.76 | 99.97 | 0.05 | 0.56 |
| CLIP ViT-L/14 | 0.77 | 0.88 | 99.97 | 0.03 | 0.75 |
| CLIP ViT-L/14@336px | 0.90 | 1.15 | 99.99 | 0.00 | 0.20 |
| SigLIP ViT-B/16 | 0.86 | 1.06 | 99.99 | 0.01 | 0.08 |
4.5.4 Analysis of the Backbone Visual-Language Model
We evaluate the effectiveness of our method when integrated with different backbone vision-language models, including CLIP ViT-B/16, CLIP ViT-L/14, CLIP ViT-L/14@336px, and SigLIP ViT-B/16. The corresponding results are presented in Table VII. Our approach transfers the general capabilities of vision-language models to enhance the performance of 3D mask presentation attack detection. From the results obtained using CLIP ViT-B/16, CLIP ViT-L/14, and SigLIP ViT-B/16, we observe a consistent trend: more powerful vision-language models lead to improved detection performance. Notably, CLIP ViT-L/14 and CLIP ViT-L/14@336px share the same model architecture but differ in input image resolution. A comparison between these two variants reveals that increasing the input resolution does not yield performance gains in our setting. This phenomenon may be attributed to the characteristics of 3D mask presentation attack detection datasets, in which face images are typically pre-aligned to a resolution of . Simply upsampling to does not introduce additional spoof patterns or enhance fine-grained details, thus offering limited benefit to the model.
@@ 4.5.5 Robustness Analysis of the Number of Entities
Entities play a pivotal role in knowledge-based prompt learning. To further investigate their impact, we analyze how the quantity of entities influences the performance of 3D mask presentation attack detection. Figure 14 presents the model performance under the HiFiMask 3DMAD protocol after randomly removing one entity from the knowledge graph, repeated over ten independent trials. Notably, when an entity is removed, all descriptions associated with its corresponding triples are also excluded. Although removing different entities results in slight performance degradation, the overall fluctuation across the five evaluation metrics remains minimal, indicating that the proposed method is robust to the removal of a single entity. This robustness can be attributed to the model’s ability to leverage fine-grained discriminative descriptions, which draw on prior knowledge extracted from large-scale language models. Such capability enables the model to effectively expand the removed entities. As a result, the information loss incurred by removing a single entity can often be compensated by the contextual expansion of related entities, thereby leading to negligible performance variation.
To further evaluate the model sensitivity to entity quantity, we perform pairwise differential t-tests on the performance distributions obtained under four conditions: randomly removing one, three, six, and ten entities, respectively, under the HiFiMask 3DMAD protocol. Figure 15 summarizes the results. The box plots illustrate the distribution of HTER scores across repeated trials for each level of entity removal. As the number of removed entities increases, we observe a gradual decline in median performance, alongside an expansion in both the interquartile range and whisker lengths, indicating increased variability. Pairwise statistical significance results are annotated above the plots, where ”ns” denotes no statistically significant difference, and an increasing number of asterisks (*) indicates stronger significance. It is evident that the performance difference between removing one and three entities is marginal, whereas the performance degradation becomes statistically significant when six or more entities are removed. In addition, the ribbon error line further visualizes the HTER variation trend across removal levels. The central line, representing the mean HTER, increases steadily as more entities are removed, while the ribbon width grows accordingly, reflecting growing performance variability. Overall, the performance of the model deteriorates with the incremental removal of entities, particularly when the number of removals increases beyond the point where compensation through remaining entities is no longer effective. These findings underscore the critical role of knowledge graph entities in enhancing the performance of 3D mask presentation attack detection.
4.6 Generalization to Diverse Attack Types.
| Method | CSW | Method | CH | ||
| HTER | AUC | HTER | AUC | ||
| ViT [63] | 21.04 | 89.12 | ConvMLP [64] | 48.02 | 53.45 |
| CLIP-V [65] | 20.00 | 87.72 | PipeNet [66] | 49.94 | 51.60 |
| CLIP [65] | 17.05 | 89.37 | FeatherNet [67] | 57.70 | 38.54 |
| FGPL [68] | 14.05 | 92.65 | FlexModal [69] | 57.70 | 66.17 |
| CoOp [20] | 9.52 | 90.49 | FaceBagNet [70] | 47.79 | 53.47 |
| CFPL [47] | 9.04 | 96.48 | ViT-S/16 [71] | 49.00 | 52.33 |
| S-CPTL [48] | 8.99 | 94.01 | mmFAS [72] | 36.92 | 69.85 |
| Ours | 7.76 | 98.03 | Ours | 15.43 | 92.68 |
In addition to evaluating performance on datasets containing only 3D mask attacks, we further design open-set scenarios involving diverse physical attack types and digital adversarial attacks to assess the generalization capability of our method.
Generalization to diverse physical attack types. The evaluation experiments on diverse physical attack types are conducted based on the CASIA-SURF (S) [73], CASIA-SURF CeFA (C in short) [74], WMCA (W in short) [36], and HQ-WMCA (H in short) [75] datasets. The WMCA and HQ-WMCA datasets include 7 and 10 different attack types, respectively, such as print attacks, replay attacks, mannequin heads, paper mask, rigid mask, flexible mask, glasses, wigs, tattoo, and makeup. We construct the CS W protocol as S-CPTL [48], using the attack-rich WMCA dataset as the test set and the CASIA-SURF and CASIA-SURF CeFA datasets as the training data. Furthermore, we introduce a more challenging C H protocol as mmFAS [72], which reduces the number of datasets in the training data while increasing the diversity of attack types in the test data, to further verify the generalization ability of our method in more universal scenarios.
The experimental results are presented in Table VIII. As shown, our method outperforms existing SOTA methods on the HTER metric by 13.68% and 58.20% under the two protocols. This substantial improvement demonstrates the strong generalization capability of our method to diverse attack types. On the left of Table VIII, the comparison methods are mostly prompt learning-based approaches, all using the same backbone vision-language model, CLIP ViT-B/16, as ours. The significant gains under the CS W protocol further highlight the effectiveness of our knowledge graph prompts. On the right of Table VIII, the evaluated methods are primarily multi-modal fusion approaches. Notably, our method leverages only visible-light modality images for both training and testing, yet still achieves considerable improvements under the C H protocol. This demonstrates that our approach can effectively generalize to cross-dataset, diverse attack types using only a single modality and a single training dataset.
| Method | HTER | EER | AUC | B@A=0.1 | B@A=0.01 |
| FASTEN [12] | 44.71 | 47.06 | 52.25 | 97.06 | 100.00 |
| FLIP [22] | 4.12 | 2.21 | 99.99 | 0.00 | 0.00 |
| Ours | 2.65 | 7.15 | 99.99 | 0.00 | 1.76 |
Generalization to digital adversarial attacks. The evaluation experiments on digital adversarial attacks are conducted based on the HiFiMask and 3DMAD datasets. We use the HiFiMask dataset as the training set. For the test set, we construct a combination of real face images from the 3DMAD dataset and adversarial attack samples generated by AT3D [76] using the real face images from the 3DMAD dataset as original images. The experimental results are presented in Table IX. The performance of the FASTEN method is obtained using the officially released pretrained model. As shown, FASTEN exhibits poor generalization to adversarial samples. In contrast, prompt learning-based methods demonstrate significantly better generalization ability to adversarial attacks. This indicates that leveraging prompt learning to tap into the general knowledge of vision-language models is beneficial for handling adversarial attacks. Compared to the simple category name prompts of FLIP, our method achieves a 37.16% improvement on the HTER metric, further validating the effectiveness of the proposed knowledge graph prompts.
| Method | Inference Times | Training Parameters | Frozen Parameters |
| FASTEN [12] | 23.06 ms | 39.8M | 162.5M |
| Ours | 3.91 ms | 12.7M | 149.6M |
4.7 Computational Complexity Analysis.
We compare the computational complexity of our method with the baseline method FASTEN in terms of inference time and model parameters. The experimental results are presented in Table X. Our method is single-frame based, with an average inference time of 3.91 ms, while FASTEN is a multi-frame based method, with an average inference time of 23.06 ms. We freeze the CLIP backbone (149.6M parameters) during training and optimize only lightweight prompt parameters and filters, totaling 12.7M parameters. In contrast, FASTEN involves multi-frame processing and spatiotemporal modeling, requiring training a FlowNetS (38M parameters) and a MobileNetV3-Small (1.8M parameters) backbone, along with a pre-trained FlowNet2.0 (162.5M parameters) used for generating ground-truth optical flow labels prior to training. All these results indicate that our method holds an advantage in terms of computational complexity.